OpenStudy (ilikestudying140898):

How many terms are in the expansion (x + y)10. 10 13 12 11 I choose 10

TheSmartOne (thesmartone):

The question is \(\sf (x+y)^{10}\) right?

OpenStudy (ilikestudying140898):

@TheSmartOne yes

jimthompson5910 (jim_thompson5910):

It might help to look at smaller expansions \[\Large (x+y)^2 = x^2+2xy+y^2 \ \rightarrow \ \text{3 terms}\] \[\Large (x+y)^3 = x^3+3x^2y+3xy^2+y^3 \ \rightarrow \ \text{4 terms}\] and see what pattern you can find

TheSmartOne (thesmartone):

^^

TheSmartOne (thesmartone):

After seeing what Jim posted, what do you think the correct answer will be? :)

OpenStudy (ilikestudying140898):

10.

TheSmartOne (thesmartone):

You didn't fully understand what Jim posted... \(\sf \Large (x+y)^\color{red}{2} \to \color{red}{3}~ terms\) \(\sf \Large (x+y)^\color{red}{3} \to \color{red}{4}~ terms\) \(\sf \Large (x+y)^\color{red}{4} \to \color{red}{5}~ terms\) \(\sf \Large (x+y)^\color{red}{5} \to \color{red}{6}~ terms\) \(\sf \Large (x+y)^\color{red}{6} \to \color{red}{7}~ terms\) ...

OpenStudy (ilikestudying140898):

(x+y)^10 is 10 terms

jimthompson5910 (jim_thompson5910):

@TheSmartOne is basically saying that for any nonnegative integer n \[\Large (x+y)^n \rightarrow \text{n+1 terms}\]

jimthompson5910 (jim_thompson5910):

for instance, if n = 1, then \[\Large (x+y)^n \rightarrow \text{n+1 terms}\] \[\Large (x+y)^1 \rightarrow \text{1+1 terms}\] \[\Large (x+y)^1 = x+y \rightarrow \text{2 terms}\]

TheSmartOne (thesmartone):

that has been our whole point, so that way you would revise your answer. :)

OpenStudy (ilikestudying140898):

thank yu

OpenStudy (ilikestudying140898):

jim also

TheSmartOne (thesmartone):

no problem

jimthompson5910 (jim_thompson5910):

you're welcome